Derivative f(x)=x^3×e^x
1
Expert's answer
2021-09-27T04:20:52-0400

Let us find the derivative of function "f(x)=x^3e^x" using the product rule "(f(x)g(x))'=f'(x)g(x)+f(x)g'(x):"


"f'(x)=3x^2e^x+x^3e^x=x^2(3+x)e^x."